# PairSem: LLM-Guided Pairwise Semantic Matching for Scientific Document Retrieval

### v1

@@ 1. Introduction
Dense retrieval (Karpukhin et al., 2020; Xiong et al., 2020; Wang et al., 2022) has recently emerged as the predominant approach for document search. These models, often built on pre-trained language models, are fine-tuned to map queries and documents into a shared vector space, where relevance is measured by embedding similarity. While effective in general web search (Karpukhin et al., 2020; Singh et al., 2023; Izacard et al., 2021), their performance often falls short in scientific domains (Cai et al., 2024; Kang et al., 2025). One limitation stems from pre-training on general-domain text, which provides limited understanding of specialized terminology and domain-specific user intent (Jin et al., 2023; Xu et al., 2024). Another challenge is that dense embeddings provide only holistic document-level representations, making it difficult to capture fine-grained details such as the target property of a specific chemical compound (Weller et al., 2025; Shavarani and Sarkar, 2025; Chen et al., 2022). In addition, constructing the annotated query–document pairs required for fine-tuning is costly and relies heavily on domain expertise, limiting adaptability to specialized fields (Kang et al., 2024a; Zhang et al., 2025a).
@@
Recent advances in LLMs (Hurst et al., 2024; Anthropic, 2024; Grattafiori et al., 2024) have encouraged new approaches that generate rich semantics relevant to queries and documents, by using the powerful comprehension ability of LLMs. Early methods prompted LLMs to generate pseudo queries (Gao et al., 2023), keywords (Mackie et al., 2023; Boudin et al., 2020), or related sentences (Zheng et al., 2020; Lei et al., 2024) to expand the original texts. However, these approaches still encode the generated semantics together with the original texts into a single embedding, which limits the ability to capture fine-grained entities such as "perfluorinated acid" or "metal-organic framework" in Figure 1. To address this, recent work (Kang et al., 2024a; Zhang et al., 2025a) builds a separate list of semantic entities (e.g., yellow box in Figure 1), explicitly linking documents with scientific concepts drawn from the pre-defined set (Shen et al., 2018). Rather than integrating the semantic entities into document embeddings, these methods incorporate an additional semantic matching score alongside the embedding-based score, enhancing retrieval in scientific domains.
@@
While this semantic matching improves alignment between queries and documents, most existing methods treat scientific concepts as independent fragments. This approach overlooks the multi-faceted characteristics of a scientific entity, where meaning often emerges from the interplay between entities and their associated aspects. For instance, in Figure 1, both documents mention "metal-organic framework", sharing most of semantic entities (yellow box). However, Document A describes "active site" (aspect) of "metal-organic framework" (entity), while Document B highlights "structural stability" (another aspect) of the same entity. By collapsing these diverse entity–aspect relationships into a single-level flat list, current methods struggle to capture fine-grained matches between queries and documents. Indeed, our analysis demonstrates that entities are frequently associated with multiple aspects (6.75 on average in ChemLit-QA dataset). This motivates the need for a new retrieval framework that explicitly models entities together with their multiple aspects, enabling more precise and context-aware matching in scientific document retrieval.
@@ 3. Related Work
Pre-trained dense retrievers. Pre-trained dense retrievers map texts into continuous Euclidean spaces, where semantic relevance between queries and documents is captured through vector similarity. This paradigm has become fundamental in modern information retrieval (IR) (Faloutsos and Oard, 1998; Kobayashi and Takeda, 2000; Mitra and Chaudhuri, 2000) because of its state-of-the-art performance (Zhao et al., 2024; Hambarde and Proença, 2023). Early studies (Karpukhin et al., 2020; Ren et al., 2021; Xiong et al., 2020; Luan et al., 2021; Zhan et al., 2021; Qu et al., 2021; Izacard et al., 2021) built upon general-domain pre-trained transformer models (Vaswani et al., 2017) such as BERT (Devlin et al., 2019) and ERNIE (Zhang et al., 2019), and demonstrated the effectiveness of dense retrieval by fine-tuning these models on query-document pairs using various negative sampling strategies. Subsequent works, such as SciBERT (Beltagy et al., 2019), SPECTER (Singh et al., 2023), and BMRetriever (Xu et al., 2024), focused on scientific retrieval by fine-tuning retriever models on domain-specific data. Recently, retrievers built upon large language models (LLMs) (Hurst et al., 2024; Yang et al., 2025b; Grattafiori et al., 2024), such as GritLM (Muennighoff et al., 2025), NV-embed (Lee et al., 2025), EmbeddingGemma (Vera et al., 2025), and Qwen3-Embedding (Yang et al., 2025b), have further leveraged their strong text understanding capabilities, substantially improving the performance of dense retrieval.
@@
LLM-enhanced retrieval. The emergence of LLMs (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023) brought substantial advances to information retrieval. The pretrained knowledge and instruction-following capabilities of LLMs can improve retrieval effectiveness without additional query–document relevance annotations.
@@ 4.1.1. Zero-shot pair generation
Additionally, we leverage the ability of LLMs for the structured output (Liu et al., 2024; Yang et al., 2025a), with a format prompt .
@@ 4.1.2. Corpus-based entity/aspect set construction
However, we cannot naively merge those concepts in the same cluster to the same entity/aspect name. For example, "1-d structure" and "3-d structure" are closely located in the embedding space, but they indeed represent distinct concepts that cannot be merged into a single aspect. To tackle this, we exploit LLMs’ strong entity comprehension ability (Viswanathan et al., 2024; Zhang et al., [n. d.]). This allows us to readily integrate different surface names for the same concept, or distinguish distinct concepts under similar semantics. Specifically, for each entity cluster , we prompt LLMs to detect a set of synonyms and merge them by generating a representative entity .
@@ 5.1.3. Base retrievers
We employ two representative dense retrievers: SPECTER2 (Singh et al., 2023) and Contriever-MS (Izacard et al., 2021). SPECTER2 is trained on scientific metadata (e.g., authors, citations), while Contriever-MS is fine-tuned on the MS MARCO dataset (Nguyen et al., 2016), which contains (q,d) pairs from general domains. Additionally, we adopt Qwen3-Embedding-8B (Yang et al., 2025b), an LLM-based embedding model that achieves state-of-the-art performance across diverse retrieval tasks. All base retrievers are sourced from their official Hugging Face repositories (Wolf et al., 2019).
@@ 5.3. Study of PairSem
Supplementary results, including the evaluation of entity/aspect predictors and a case study, are presented in Appendix A.2.
@@ 5.3.1. Analysis on pair generation
Table 2 reports statistics on pair generation with GPT-4.1-mini for the ChemLit-QA corpus. For each step in §4.1, we present the time required for LLM generation (Time), the corresponding API costs (Cost), and the output statistics (Output). From 1,528 documents, PairSem constructs sets of 3,352 entities (merged from 5,032) and 2,813 aspects (merged from 4,649), and generates 21.86 pairs per document on average. It is worth noting that the average number of generated pairs is substantially increased when the pair generation is augmented with candidate entities and aspects (13.55 21.86). This finding shows that candidate-augmented pair generation addresses two key challenges in zero-shot pair generation: it grounds the pairs in synonym-merged sets and also prevents LLMs from missing relevant pairs. Figure 4 further presents the histogram of the number of unique aspects linked to each entity in the corpus. Among the 3,352 entities, more than half are associated with more than two aspects, and the average number of unique aspects per entity is 6.75. This supports our motivation for generating pairwise semantics rather than relying on flat entity sets.
@@ 5.3.3. Results with domain-specific retrievers
While our main objective is to enhance general retrievers for specific target domains, we also demonstrate that PairSem can improve domain-specific retrievers without relying on human annotations. We select two recent biomedical retrievers, MedCPT (Jin et al., 2023) and BMRetriever-1B (Xu et al., 2024), where domain-specific advancements have been actively explored. MedCPT is trained on large-scale search logs and papers from PubMed, and BMRetriever-1B is trained on biomedical corpora such as bioRxiv888https://www.biorxiv.org/ and multiple medical QA datasets (McCreery et al., 2020; Herlihy and Rudinger, 2021; Pal et al., 2022).
Table 4 presents the retrieval performance of these retrievers and our enhancements on the SciFact dataset. We find that PairSem still improves their performance even though both retrievers are already fine-tuned on domain-specific data. In biomedical contexts, capturing fine-grained entities (e.g., "homocysteine") and linking them with specific aspects (e.g., "blood level") is critical. Despite domain-specific fine-tuning, it still remains difficult for holistic document-level embeddings to capture such detailed and structured semantics (Kang et al., 2024a; Shavarani and Sarkar, 2025). By generating pairwise semantics and explicitly leveraging them for matching, PairSem successfully enhances the effectiveness of these domain-specific retrievers.

### v2

@@ 1. Introduction
Dense retrieval (Karpukhin et al., 2020; Xiong et al., 2020; Wang et al., 2022) has recently emerged as the predominant approach for document search. These models, often built on pre-trained language models, are fine-tuned to map queries and documents into a shared vector space, where relevance is measured by embedding similarity. While effective in general web search (Karpukhin et al., 2020; Singh et al., 2023; Izacard et al., 2021), their performance often falls short in scientific domains (Cai et al., 2024; Kang et al., 2025). One limitation stems from pre-training on general-domain text, which provides limited understanding of specialized terminology and domain-specific user intent (Jin et al., 2023; Xu et al., 2024). Another challenge is that dense embeddings provide only holistic document-level representations, making it difficult to capture fine-grained details such as the target property of a specific chemical compound (Weller et al., 2025; Shavarani and Sarkar, 2025; Chen et al., 2022). In addition, constructing the annotated query–document pairs required for fine-tuning is costly and relies heavily on domain expertise, limiting adaptability to specialized fields (Kang et al., 2024a; Zhang et al., 2025a; Shi et al., 2025; Lee et al., 2026).
@@
Recent advances in LLMs (Hurst et al., 2024; Anthropic, 2024; Grattafiori et al., 2024) have encouraged new approaches that generate rich semantics relevant to queries and documents, by using the powerful comprehension ability of LLMs. Early methods prompted LLMs to generate pseudo queries (Gao et al., 2023), keywords (Mackie et al., 2023; Boudin et al., 2020), or related sentences (Zheng et al., 2020; Lei et al., 2024) to expand the original texts. However, these approaches still encode the generated semantics together with the original texts into a single embedding, which limits the ability to capture fine-grained entities such as "nitro group" or "formyl group" in Figure 1. To address this, recent work (Kang et al., 2024a; Zhang et al., 2025a; Kweon et al., 2025b) builds a separate list of semantic entities (e.g., yellow box in Figure 1), explicitly linking documents with scientific concepts drawn from the pre-defined set (Shen et al., 2018). Rather than integrating the semantic entities into document embeddings, these methods incorporate an additional semantic matching score alongside the embedding-based score, enhancing retrieval in scientific domains.
@@
While this semantic matching improves alignment between queries and documents, most existing methods treat scientific concepts as independent fragments. This approach overlooks the multi-faceted characteristics of a scientific entity, where meaning often emerges from the interplay between entities and their associated aspects. For instance, in Figure 1, both documents mention "nitro group", sharing most of semantic entities (yellow box). However, Document A describes "electron accepting ability" (aspect) of "nitro group" (entity), while Document B highlights "C4 substituent" (another aspect) of the same entity. By collapsing these diverse entity–aspect relationships into a single-level flat list, current methods struggle to capture fine-grained matches between queries and documents. Indeed, our analysis demonstrates that entities are frequently associated with multiple aspects (6.75 on average in ChemLit-QA dataset). This motivates the need for a new retrieval framework that explicitly models entities together with their multiple aspects, enabling more precise and context-aware matching in scientific document retrieval.
@@ 3. Related Work
Pre-trained dense retrievers. Pre-trained dense retrievers map texts into continuous Euclidean spaces, where semantic relevance between queries and documents is captured through vector similarity. This paradigm has become fundamental in modern information retrieval (IR) (Faloutsos and Oard, 1998; Kobayashi and Takeda, 2000; Mitra and Chaudhuri, 2000) because of its state-of-the-art performance (Zhao et al., 2024; Hambarde and Proença, 2023). Early studies (Karpukhin et al., 2020; Ren et al., 2021; Xiong et al., 2020; Luan et al., 2021; Zhan et al., 2021; Qu et al., 2021; Izacard et al., 2021) built upon general-domain pre-trained transformer models (Vaswani et al., 2017) such as BERT (Devlin et al., 2019) and ERNIE (Zhang et al., 2019), and demonstrated the effectiveness of dense retrieval by fine-tuning these models on query-document pairs using various negative sampling strategies. Subsequent works, such as SciBERT (Beltagy et al., 2019), SPECTER (Singh et al., 2023), and BMRetriever (Xu et al., 2024), focused on scientific retrieval by fine-tuning retriever models on domain-specific data. Recently, retrievers built upon large language models (LLMs) (Hurst et al., 2024; Yang et al., 2025a; Grattafiori et al., 2024), such as GritLM (Muennighoff et al., 2025), NV-embed (Lee et al., 2025), EmbeddingGemma (Vera et al., 2025), and Qwen3-Embedding (Yang et al., 2025a), have further leveraged their strong text understanding capabilities, substantially improving the performance of dense retrieval.
@@
LLM-enhanced retrieval. The emergence of LLMs (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023) brought substantial advances to information retrieval. The pretrained knowledge and instruction-following capabilities of LLMs can improve retrieval effectiveness without additional query–document relevance annotations (Kweon et al., 2025a).
@@ 4.1.1. Zero-shot pair generation
Additionally, we leverage the ability of LLMs for the structured output (Liu et al., 2024; Yang et al., 2025b), with a format prompt .
@@ 4.1.2. Corpus-based entity/aspect set construction
However, we cannot naively merge those concepts in the same cluster to the same entity/aspect name. For example, "1-d structure" and "3-d structure" are closely located in the embedding space, but they indeed represent distinct concepts that cannot be merged into a single aspect. To tackle this, we exploit LLMs’ strong entity comprehension ability (Viswanathan et al., 2024; Zhang et al., ). This allows us to readily integrate different surface names for the same concept, or distinguish distinct concepts under similar semantics. Specifically, for each entity cluster , we prompt LLMs to detect a set of synonyms and merge them by generating a representative entity .
@@ 5.1.3. Base retrievers
We employ two representative dense retrievers: SPECTER2 (Singh et al., 2023) and Contriever-MS (Izacard et al., 2021). SPECTER2 is trained on scientific metadata (e.g., authors, citations), while Contriever-MS is fine-tuned on the MS MARCO dataset (Nguyen et al., 2016), which contains (q,d) pairs from general domains. Additionally, we adopt Qwen3-Embedding-8B (Yang et al., 2025a), an LLM-based embedding model that achieves state-of-the-art performance across diverse retrieval tasks. All base retrievers are sourced from their official Hugging Face repositories (Wolf et al., 2019).
@@ 5.3. Study of PairSem
Supplementary results (e.g., evaluation of entity/aspect predictors and hyperparameter study) are presented in Appendix A.2.
@@ 5.3.1. Analysis on pair generation
Table 2 reports statistics on pair generation for the ChemLit-QA corpus. For each step in §4.1, we present the time required for LLM generation (Time), the corresponding API costs (Cost), and the output statistics (Output). From 1,528 documents, PairSem constructs sets of 3,352 entities (merged from 5,032) and 2,813 aspects (merged from 4,649), and generates 21.86 pairs per document on average. It is worth noting that the average number of generated pairs is substantially increased when the pair generation is augmented with candidate entities and aspects (13.55 21.86). Figure 4 further presents the histogram of the number of unique aspects linked to each entity in the corpus. Among the 3,352 entities, more than half are associated with more than two aspects, and the average number of unique aspects per entity is 6.75. This supports our motivation for generating pairwise semantics rather than relying on flat entity sets.
@@ 5.3.3. Results with domain-specific retrievers
While our main objective is to enhance general retrievers for specific target domains, we also demonstrate that PairSem can improve domain-specific retrievers without relying on human annotations. We select two recent biomedical retrievers, MedCPT (Jin et al., 2023) and BMRetriever-1B (Xu et al., 2024). MedCPT is trained on large-scale search logs and papers from PubMed, and BMRetriever-1B is trained on biomedical corpora such as bioRxiv888https://www.biorxiv.org/ and multiple medical QA datasets (McCreery et al., 2020; Herlihy and Rudinger, 2021; Pal et al., 2022).
Table 4 shows that PairSem improves their performance even though both retrievers are already fine-tuned on domain-specific data. In biomedical contexts, capturing entities (e.g., "homocysteine") and linking them with specific aspects (e.g., "blood level") is critical. Despite domain-specific fine-tuning, it still remains difficult for holistic document-level embeddings to capture such detailed and structured semantics (Kang et al., 2024a; Shavarani and Sarkar, 2025). By generating pairwise semantics and explicitly leveraging them for matching, PairSem successfully enhances the effectiveness of these domain-specific retrievers.
@@ 6. Conclusion
Acknowledgements.
This work was supported by Samsung Research Funding & Incubation Center of Samsung Electronics (SRFC-IT2402-05, South Korea), by MSIT (No. RS-2024-00335873, South Korea), and by Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF (No. 2019897, United States).References
- Gpt-4 technical report. arXiv preprint arXiv:2303.08774. Cited by: §5.1.4.
- LitSearch: a retrieval benchmark for scientific literature search. In EMNLP, Cited by: §5.1.1.
- Claude 3 model family. Note: https://assets.anthropic.com/m/61e7d27f8c8f5919/original/Claude-3-Model-Card.pdfAccessed: 2025-05-06 Cited by: §1.
- SciBERT: a pretrained language model for scientific text. In EMNLP, Cited by: §3.
- Keyphrase generation for scientific document retrieval. In ACL, Cited by: §1, §3.
- Language models are few-shot learners. In NeurIPS, Cited by: §3.
- MixGR: enhancing retriever generalization for scientific domain through complementary granularity. In EMNLP, Cited by: §1, §1, §3.
- Salient phrase aware dense retrieval: can a dense retriever imitate a sparse one?. In Findings of EMNLP, Cited by: §1, §3.
- Palm: scaling language modeling with pathways. Journal of Machine Learning Research 24 (240), pp. 1–113. Cited by: §3.
- Reciprocal rank fusion outperforms condorcet and individual rank learning methods. In SIGIR, Cited by: §3, §4.2.2.
- BERT: pre-training of deep bidirectional transformers for language understanding. In NAACL, Cited by: §3.
- The faiss library. arXiv preprint arXiv:2401.08281. Cited by: §2.
- A survey of information retrieval and filtering methods. Cited by: §3.
- Precise zero-shot dense retrieval without relevance labels. In ACL, Cited by: §A.1.2, §1, §3, Table 1, Table 1, Table 1, 1st item.
- The llama 3 herd of models. arXiv preprint arXiv:2407.21783. Cited by: §1, §3.
- Information retrieval: recent advances and beyond. IEEE Access 11 (), pp. 76581–76604. External Links: Document Cited by: §3.
- MedNLI is not immune: natural language inference artifacts in the clinical domain. In ACL, Cited by: §5.3.3.
- Gpt-4o system card. arXiv preprint arXiv:2410.21276. Cited by: §1, §3.
- Unsupervised dense information retrieval with contrastive learning. arXiv preprint arXiv:2112.09118. Cited by: §1, §3, §5.1.3.
- Cumulated gain-based evaluation of ir techniques. ACM Transactions on Information Systems (TOIS) 20 (4), pp. 422–446. Cited by: §5.1.2.
- Instruct and extract: instruction tuning for on-demand information extraction. In EMNLP, Cited by: §3.
- Medcpt: contrastive pre-trained transformers with large-scale pubmed search logs for zero-shot biomedical information retrieval. Bioinformatics 39 (11), pp. btad651. Cited by: §1, §5.3.3.
- Improving retrieval in theme-specific applications using a corpus topical taxonomy. In WWW, Cited by: §A.1.2, §1, §1, §1, §3, Table 1, Table 1, Table 1, 3rd item, §5.3.3.
- Improving scientific document retrieval with concept coverage-based query set generation. In WSDM, Cited by: §1, §1.
- Taxonomy-guided semantic indexing for academic paper search. In EMNLP, Cited by: §3.
- MEGClass: extremely weakly supervised text classification via mutually-enhancing text granularities. In Findings of EMNLP, Cited by: §3.
- Dense passage retrieval for open-domain question answering. In EMNLP, Cited by: §1, §3.
- Information retrieval on the web. ACM Comput. Surv. 32 (2), pp. 144–173. External Links: ISSN 0360-0300 Cited by: §3.
- Lexically-accelerated dense retrieval. In SIGIR, Cited by: §A.1.2, Table 1, Table 1, Table 1, 2nd item.
- Uncertainty quantification and decomposition for llm-based recommendation. In WWW, Cited by: §3.
- Topic coverage-based demonstration retrieval for in-context learning. In EMNLP, Cited by: §1.
- NV-embed: improved techniques for training llms as generalist embedding models. In ICLR, Cited by: §3.
- Taxocom: topic taxonomy completion with hierarchical discovery of novel topic clusters. In WWW, Cited by: §4.1.4.
- Improving scientific document retrieval with academic concept index. arXiv preprint arXiv:2601.00567. Cited by: §1.
- Corpus-steered query expansion with large language models. In EACL, Cited by: §1, §3, Table 1, Table 1, Table 1, 3rd item.
- ”We need structured output”: towards user-centered constraints on large language model output. In Extended Abstracts of the CHI, Cited by: §4.1.1.
- Sparse, dense, and attentional representations for text retrieval. Transactions of the Association for Computational Linguistics 9, pp. 329–345. Cited by: §3.
- Query rewriting in retrieval-augmented large language models. In EMNLP, Cited by: §3.
- Generative relevance feedback with large language models. In SIGIR, Cited by: §A.1.2, §1, §3, Table 1, Table 1, Table 1, 2nd item.
- Effective transfer learning for identifying similar questions: matching user questions to covid-19 faqs. In KDD, Cited by: §5.3.3.
- Weakly-supervised neural text classification. In CIKM, Cited by: §3.
- Information retrieval from documents: a survey. Information retrieval 2 (2), pp. 141–163. Cited by: §3.
- Generative representational instruction tuning. In ICLR, Cited by: §3.
- MS-marco: a human-generated machine reading comprehension dataset. Cited by: §5.1.3.
- A survey on open information extraction. In COLING, Cited by: §3.
- Structured chemistry reasoning with large language models. In ICML, Cited by: §1.
- Medmcqa: a large-scale multi-subject multi-choice dataset for medical domain question answering. In CHIL, Cited by: §5.3.3.
- Scikit-learn: machine learning in python. the Journal of machine Learning research 12, pp. 2825–2830. Cited by: §A.1.3.
- ADELIE: aligning large language models on information extraction. In EMNLP, Cited by: §3.
- RocketQA: an optimized training approach to dense passage retrieval for open-domain question answering. In NAACL, Cited by: §3.
- PAIR: leveraging passage-centric similarity relation for improving dense passage retrieval. In Findings of ACL, Cited by: §3.
- Entity retrieval for answering entity-centric questions. In Workshop on Knowledge-Augmented Methods for Natural Language Processing, NAACL, Cited by: §1, §3, §5.3.3.
- A web-scale system for scientific knowledge exploration. In ACL, Cited by: §1, §3.
- Hypercube-rag: hypercube-based retrieval-augmented generation for in-domain scientific question-answering. arXiv preprint arXiv:2505.19288. Cited by: §1.
- SciRepEval: a multi-format benchmark for scientific document representations. In EMNLP, Cited by: §1, §3, §5.1.3.
- Is ChatGPT good at search? investigating large language models as re-ranking agents. In EMNLP, Cited by: §3.
- Chemagent: self-updating memories in large language models improves chemical reasoning. In ICLR, Cited by: §1.
- LLM-based compact reranking with document features for scientific retrieval. arXiv preprint arXiv:2505.13757. Cited by: §3, §3.
- Llama: open and efficient foundation language models. arXiv preprint arXiv:2302.13971. Cited by: §3.
- Attention is all you need. In NeurIPS, Cited by: §3.
- EmbeddingGemma: powerful and lightweight text representations. arXiv preprint arXiv:2509.20354. Cited by: §3.
- Large language models enable few-shot clustering. Transactions of the Association for Computational Linguistics 12, pp. 321–333. Cited by: §4.1.2.
- Fact or fiction: verifying scientific claims. In EMNLP, Cited by: §5.1.1.
- Microsoft academic graph: when experts are not enough. Quantitative Science Studies 1 (1), pp. 396–413. Cited by: §A.1.2.
- Text embeddings by weakly-supervised contrastive pre-training. arXiv preprint arXiv:2212.03533. Cited by: §1.
- Hierarchical grouping to optimize an objective function. Journal of the American statistical association 58 (301), pp. 236–244. Cited by: §4.1.2.
- ChemLit-qa: a human evaluated dataset for chemistry rag tasks. In Workshop on AI for Accelerated Materials Design @ NeurIPS, Cited by: §5.1.1.
- On the theoretical limitations of embedding-based retrieval. arXiv preprint arXiv:2508.21038. Cited by: §1.
- Ontology-based information extraction: an introduction and a survey of current approaches. Journal of Information Science 36 (3), pp. 306–323. Cited by: §3.
- Huggingface’s transformers: state-of-the-art natural language processing. arXiv preprint arXiv:1910.03771. Cited by: §5.1.3.
- Approximate nearest neighbor negative contrastive learning for dense text retrieval. In ICLR, Cited by: §1, §3.
- BMRetriever: tuning large language models as better biomedical text retrievers. In EMNLP, Cited by: §1, §3, §5.1.2, §5.3.3.
- Zero-shot open-schema entity structure discovery. arXiv preprint arXiv:2506.04458. Cited by: §3.
- Qwen3 technical report. arXiv preprint arXiv:2505.09388. Cited by: §3, §5.1.3.
- StructEval: benchmarking llms’ capabilities to generate structural outputs. arXiv preprint arXiv:2505.20139. Cited by: §4.1.1.
- Generate rather than retrieve: large language models are strong context generators. In ICLR, Cited by: §3.
- Optimizing dense retrieval model training with hard negatives. In SIGIR, Cited by: §3.
- Scientific paper retrieval with llm-guided semantic-based ranking. In EMNLP, Cited by: §A.1.2, Table 7, Table 7, Table 7, Table 7, §1, §1, §1, §3, §3, Table 1, Table 1, Table 1, 4th item, §5.1.2.
- TELEClass: taxonomy enrichment and llm-enhanced hierarchical text classification with minimal supervision. In WWW, Cited by: §3.
- [80] ClusterLLM: large language models as a guide for text clustering. In EMNLP, Cited by: §4.1.2.
- ERNIE: enhanced language representation with informative entities. In ACL, Cited by: §3.
- Dense text retrieval based on pretrained language models: a survey. ACM Trans. Inf. Syst. 42 (4). External Links: ISSN 1046-8188 Cited by: §3.
- BERT-qe: contextualized query expansion for document re-ranking. In Findings of EMNLP, Cited by: §A.1.2, §1, Table 1, Table 1, Table 1, 1st item.
- Benchmarking retrieval-augmented generation for chemistry. In COLM, Cited by: §1.